OpenStudy (anonymous):

Clara invests $5,000 in an account that pays 6.25% interest per year. After how many years will her investment be worth $10,000?

OpenStudy (anonymous):

I think the interest is compounded yearly, if so A = $10000 P = $5000 r = 6.25% = 0.0625 A = P (1+r/100)^t where t denotes time in years 10000 = 5000 (1+0.0625)^t 10000/5000 = 1.0625^t 2 = 1.0625^t Use logarithms or a scientific calculator t = 11.5 In 11.5 years $5000 investment be worth of $10000 >>>>>>>>>>>>> In case the interest is simple interest Interest = Amount - Principal = $10000-$5000 = $5000 Time = 100*Interest / (Principal*Rate = 100*5000 / (5000*6.25) = 16 years In 16 years $5000 investment be worth of $10000 http://answers.yahoo.com/question/index?qid=20110307213635AAdjbxo